# Unified Taxonomy for Multivariate Time Series Anomaly Detection using Deep Learning

### 19 Mar 2026 v1

@@ 4.2 Output: Type, Thresholding and Granularity
where $q$ is the desired proportion of anomalies, $N$ is the number of AS values, $N_{th}$ is the number of these values that are higher than $th$ and $\hat{\gamma}$ and $\hat{\beta}$ are estimates for model parameters [OmniAnomaly, POT_and_variants]. Therefore, using this method, two parameters need to be tuned: $q$ and the threshold $th$. Alternatively, some works use ML-based algorithms for anomaly/outlier detection on top of the anomaly scores or features computed by the DL models. Some examples include the IF, OCSVM and the Local Outlier Factor (LOF) [Nguyen2021, CSL]. Finally, several works perform an exhaustive search over a predefined range of potential thresholds to identify that with the highest F1-score. In this approach, the F1-score is calculated for each candidate threshold in the range, and the optimal threshold $\delta$ is selected as the one maximizing the F1-score [USAD, GTA, Interfusion, Zhou2022_CAE-AD].
Instead of fixing a threshold for the entire inference stage, other approaches have been adapting a Time-varying Threshold. For example, the sigma-rule (14) can be used in a time-varying fashion with an Exponential Weighted Moving Average (EWMA) and an Exponential Weighted Standard Deviation (EWSD) [Matar2023], or with the Nonparametric Dynamic Thresholding (NDT) method proposed by MSL. Finally, variations of POT may also be used to adapt the threshold over time, including streaming POT (SPOT) and drift-aware SPOT (DSPOT) [POT_and_variants].
Most of the Output Granularity (Point | Subsequence) choices are Point rather than Subsequence. Even when $AS$ is computed at the window level, it is usually mapped to points — commonly the last [USAD, CTAD], but sometimes the first [ZhouX2023LAT] or middle point [WangF2023]. Point-level $AS$ can also be estimated by averaging overlapping windows [MAD-GAN, TFAD]. In contrast, Subsequence methods label entire windows directly [MSCRED, AMSL]. Although granularity is independent of evaluation, point-level approaches often use point-adjustment to align predictions with segment-level ground truth. Moreover, beyond detecting anomalies at the point or subsequence level, some studies also identify the specific metrics where anomalies occur, enabling fine-grained anomaly detection [MFGAD, STVAE].
@@ 4.4 Model: T/S dependency and Feature Extraction
Two other important dimensions of the Model part are T/S dependency and Feature Extraction. These dimensions are closely tied to the DL architectures employed within a model. Building on existing taxonomies for DL-based methods [ZamanzadehDarban2024, Choi2021], popular architectures for MTSAD were grouped into temporal architectures comprising RNN, Transformer and TCN, and spatial architectures comprising CNN, GNN, AE, VAE and GAN. Other generative architectures, such as Normalizing Flows and Diffusion Models, have also been used in MTSAD despite not directly modeling temporal or spatial dependencies. Normalizing Flows (e.g. MAF and Real NVP) transform a simple base distribution (e.g. Gaussian) into a complex target distribution via invertible mappings, allowing efficient density estimation [Kobyzev2021]. In MTSAD, anomalies are assumed to occur in low-density regions [GANF, MTGFlow]. Diffusion Models (e.g. DDPM) instead corrupt inputs with noise and learn to reverse this process for reconstruction [ImDiffusion, DDTAD]. In practice, these generative models are often combined with spatio-temporal architectures to better capture the complex MTS patterns.
@@ 4.4.4 Nested Models
In Nested models, typically, a temporal model is embedded within an AE a GAN-based model. For instance, several studies have implemented LSTM-AE architectures, where the encoder and decoder structure of an AE is composed of LSTM layers [Kieu2019, Nguyen2021]. As another example, the OmniAnomaly uses GRU incorporated into a VAE framework to create a stochastic RNN to learn latent representations and capture temporal dependencies [OmniAnomaly]. Additionally, DAEMON (adversarial Autoencoder anomaly detection interpretation) uses adversarial training in an AE-based model, where encoder and decoder networks consist of 1D-CNN, which apply filters that slide along the temporal dimension, to capture the temporal dependencies [DAEMON]. An example of a GAN-based model is MAD-GAN (Multivariate Anomaly Detection with GAN), which employs an LSTM network as the base model for both the Generator and Discriminator. MAD-GAN considers the entire dimension set simultaneously in order to capture the latent interactions among the different dimensions and uses LSTM to capture temporal dependency [MAD-GAN].
@@ 4.4.5 Parallel Models
Parallel models handle both dependencies simultaneously using a two-branch architecture, one dedicated to temporal dependencies and the other to spatial dependencies. Typically, a temporal model is employed in one branch and a spatial model in the other. For example, the STADN (AD Network using Spatial and Temporal information) models spatial dependencies using a GAT and temporal dependencies using an LSTM. The two feature representations are then concatenated and passed through stacked fully connected layers for prediction [Tian2023]. Similarly, ESTAD (Enhanced Spatio-Temporal constraints end-to-end network for AD) employs a Transformer encoder to extract temporal features in parallel with a GNN model that learns spatial dependencies. It further introduces constraint tasks in both temporal and spatial domains, such as an autoregressive temporal objective and graph contrastive learning, to enhance spatio-temporal representations [ESTAD].
@@ 5.1 Overview until 2024
Before 2019, MTSAD research was dominated by traditional ML methods, such as clustering, Support Vector Machines and tree-based models, which were valued for their interpretability and their ability to capture anomaly patterns across multiple variables. Although such works were identified in the survey, they were not included in the review since the focus lies on DL-based approaches. Readers interested in ML-based strategies are referred to existing surveys [BlzquezGarca2021, Paparrizos2025]. From 2019 onwards, a clear shift towards DL-based models can be observed. The first wave of contributions, particularly between 2019 and 2022, introduced disruptive architectural innovations, such as OmniAnomaly [OmniAnomaly], MAD-GAN [MAD-GAN], MSCRED [MSCRED], USAD [USAD], MTAD-GAT [MTAD-GAT], GDN [GDN], TranAD [TranAD], GTA [GTA] and Anomaly Transformer [AnomalyTransformer]. More recent works, published between 2023 and 2024, tend to focus on incremental improvements, either by fine-tuning existing architectures, modifying loss functions, or refining input and output representations. Nevertheless, notable exceptions remain, with models such as CAE-M [CAE-M] and DCdetector [DCdetector] introducing novel architectural concepts after 2022.
@@
The categorization of 409 articles highlights several trends in the field for 2019–2024. Most studies rely on raw time series organized into sliding windows while preserving the temporal order, although there is a growing tendency to represent time series as images to capture spatial correlations. Regression-based settings dominate, and distance-based losses, such as MSE, remain the most widely adopted, even though hybrid loss functions have been increasingly explored. In temporal/spatial dependency modeling, sequential, nested, and other complex structures prevail, with nested and heterogeneous architectures being particularly common. Moreover, parallel structures are gaining visibility due to their scalable nature, enabling simultaneous modeling of multiple dependencies and improving computational efficiency. Dedicated feature extraction modules are relatively rare, but when present, they typically focus on extracting spatio-temporal features. With respect to the output, anomaly scores are most often computed through reconstruction errors, although hybrid strategies that combine reconstruction errors with prediction errors or latent space deviations are gaining popularity. Detection continues to be primarily carried out at the point level, with thresholds that are generally optimal but time-invariant. The choice of architecture has also evolved, with Transformers and GNN becoming increasingly prominent. Transformers are particularly effective for temporal modeling, while GNN are widely adopted to capture spatial dependencies. More recent directions include the exploration of generative models, such as Normalizing Flows and Diffusion models, as well as contrastive learning strategies. The application of Large Language Models (LLM) to MTSAD is still at an early stage, but their adoption is increasing, reflecting the broader impact of Transformer-based architectures in time series research.
@@
The dimensions of the proposed taxonomy are independent, and each article is classified into a single category per dimension. Nevertheless, certain patterns may emerge across dimensions. The most frequent combination includes No Preprocessing, Window Format, Time Content, Reconstruction Type, an Optimal but Time-invariant Threshold, Point Granularity, Regression Task, ELBO Loss Function, Nested T/S Dependency, and No Feature Extraction module. However, this setup occurs in only 18 out of the 409 categorized papers, highlighting the heterogeneity of approaches. Since the Loss Function, Optimal Threshold and Time-varying Threshold dimensions highly contributed to this variability, they were ignored for subsequent analyses. When ignoring these three dimensions, the number of papers matching this combination rises to 79 out of 409 (i.e. 19.32%). This joint proportion is larger than the product of proportions over each dimension, which correponds to the case of categories being chosen independently ($\sim$5%). This indicates a positive association such that that the presence of a category in one dimension increases the likelihood of a category in the remaining dimensions.
@@ 5.2 Some developments in 2025
Input-related dimensions (Pre-processing, Format and Content) show continuity with earlier years with some broadening. Most recent contributions continue to rely on Window-based formats for modeling subsequences. As examples, MLAD adopt fixed-length sliding windows to capture localized temporal dependencies, while TRLP introduce multi-resolution pyramid windows that model dynamics at multiple temporal scales. Together, these papers exemplify how contemporary research refines rather than replaces traditional window-based representations in MTSAD. At the same time, several studies move beyond purely temporal segmentation by integrating Window+Image representations or frequency-transformed inputs. For instance, FDTAD incorporate frequency-domain features alongside temporal windows, and FreqWave-TranDuD make use of wavelet-based frequency transformations to enhance multiscale sensitivity. These trends consolidate the Preprocessing Input dimension within the taxonomy and highlight a broader shift toward hybrid representations that merge temporal, spatial and frequency-domain information.
@@
At the Output level, the categories of Reconstruction and Prediction remain central, but 2025 research expands the use of Representation-based deviations and Hybrid anomaly scores. AD2T (Association Discrepancy Dual-decoder Transformer) combines reconstruction, prediction and association discrepancy in a dual-decoder framework [AD2T], while TRLP (Tranformer-based Reconstruction-Limited Probability) introduces probabilistic scoring based on representation uncertainty [TRLP]. Furthermore, adaptive thresholding mechanisms gained traction e.g. Mixer-transformer employs time-varying thresholds to reduce false positives [Mixer-transformer], while boundary-aware ensembles generate thresholds based on estimated uncertainty [He2025]. Together, these papers reinforce the Hybrid type output category and extend the relevance of Time-varying thresholds. For Granularity, most contributions still detect anomalies at the Point level. For example, methodologies such as MLAD (Multi-task Learning Anomaly Detection) and MGDRF (Multi-Grain Dynamic Receptive Field) explicitly model multi-scale subsequences, yet they assign anomaly scores to each time point by considering its temporal context [MLAD, MDGRF]. These developments indicate that point granularity anomaly detection using window-based inputs remains a primary focus in 2025. The Model part is where 2025 contributions show the greatest methodological innovation. In the Task dimension, explicit Classification formulations remain uncommon and Regression continues to dominate, with models learning to identify deviations as anomalies (e.g. [MLAD, TRLP, DiffANT, TiTAD, TSAD]). Moreover, several 2025 approaches integrate prediction with other representation learning (e.g. contrastive, variational or adversarial) to enhance the discriminative capacity of latent features (e.g. [RGAnomaly, DyGCL, IGANomaly, AVAE]). As a consequence, the Loss Function dimension shows some relevant developments. Beyond the classical reconstruction losses (SumE, ELBO), the 2025 research incorporates more frequent Contrastive objectives (CARLA [CARLA], DMAP-DDCL [DMAP-DDCL]), Adversarial losses (EH-GAM-EGAN [EH-GAM-EGAN], Quantum GAN [Quantum_GAN]) and multi-objective criteria that merge reconstruction, prediction and representation learning (AD2T [AD2T], STAMP [STAMP]). These advances consolidate the Multi-objective category as a mainstream choice for the Loss Function dimension.

### 24 Apr 2026 v2

@@ 4.2 Output: Type, Thresholding and Granularity
where $q$ is the desired proportion of anomalies, $N$ is the number of AS values, $N_{th}$ is the number of these values that are higher than $th$ and $\hat{\gamma}$ and $\hat{\beta}$ are estimates for model parameters [OmniAnomaly, POT_and_variants]. Therefore, using this method, two parameters need to be tuned: $q$ and the threshold $th$. Alternatively, some works use ML-based algorithms for anomaly/outlier detection on top of the anomaly scores or features computed by the DL models. Some examples include the IF, OCSVM and the Local Outlier Factor (LOF) [Nguyen2021, CSL]. Finally, several works perform an exhaustive search over a predefined range of potential thresholds to identify that with the highest F1-score. In this approach, the F1-score is calculated for each candidate threshold in the range, and the optimal threshold $\delta$ is selected as the one maximizing the F1-score [USAD, GTA, Interfusion, Zhou2022_CAE-AD]. This approach provides an upper limit to the performance of the method. However, contrary to POT and ML-based approaches, it is not feasible in an online setting, where the threshold needs to be set before inference, without access to labeled anomalies.
Instead of fixing a threshold for the entire inference stage, other approaches have been adapting a Time-varying Threshold. For example, the sigma-rule (14) can be used in a time-varying fashion with an Exponential Weighted Moving Average (EWMA) and an Exponential Weighted Standard Deviation (EWSD) [Matar2023], or with the Nonparametric Dynamic Thresholding (NDT) method proposed by MSL. Finally, variations of POT may also be used to adapt the threshold over time, including streaming POT (SPOT) and drift-aware SPOT (DSPOT) [POT_and_variants]. Compared to a time-invariant threshold, a time-varying threshold has the advantage of dynamically adjusting to the data, which is useful when the data is non-stationary. However, it has the disadvantage of needing to be computed continuously and may be sensitive to noise, as fluctuations in the AS may adapt the threshold unnecessarily, reducing detection performance.
Most of the Output Granularity (Point | Subsequence) choices are Point rather than Subsequence. Even when $AS$ is computed at the window level, it is usually mapped to points — commonly the last [OmniAnomaly, USAD, CTAD], but sometimes the first [ZhouX2023LAT] or middle point [WangF2023]. Point-level $AS$ can also be estimated by averaging overlapping windows [MAD-GAN, TFAD]. In contrast, Subsequence methods label entire windows directly [MSCRED, CAE-M, AMSL]. While subsequence granularity treats the entire window as the detection unit, point granularity enables finer temporal resolution, allowing more precise localization of anomalies. Although granularity is independent of evaluation, point-level approaches often use point-adjustment to align predictions with segment-level ground truth. Moreover, beyond detecting anomalies at the point or subsequence level, some studies also identify the specific metrics where anomalies occur, enabling fine-grained anomaly detection [MFGAD, STVAE].
@@ 4.3 Model: Task and Loss Function
The choice of loss function defines the optimization criterion used to train the model. SumE and Cross-Entropy losses minimize the error directly, making them simple and computationally efficient, but they do not account for uncertainty. NLL and ELBO address this by modeling the data distribution explicitly, enabling uncertainty quantification, at the cost of requiring distributional assumptions. Moving further from direct supervision, the Adversarial loss implicitly learns normality through a minimax game, allowing the model to capture complex data distributions, but it introduces training instability. Contrastive losses take a different perspective, shifting the focus from error minimization to representation, ensuring that normal samples cluster together in latent space. Finally, Multi-objective formulations combine complementary losses to capture multiple aspects of normality, at the cost of increased complexity.
@@ 4.4 Model: T/S dependency and Feature Extraction
Two other important dimensions of the Model part are T/S dependency and Feature Extraction. Although T/S Dependency and Feature Extraction relate to the same underlying DL architectures, they capture different aspects of model design. T/S Dependency describes how the model globally organizes the modeling of temporal and spatial dependencies, while Feature Extraction indicates wheter such a dedicated module is employed to extract temporal, spatial, or spatio-temporal features.
Building on existing taxonomies for DL-based methods [ZamanzadehDarban2024, Choi2021], popular architectures for MTSAD were grouped into temporal architectures comprising RNN, Transformer and TCN, and spatial architectures comprising CNN, GNN, AE, VAE and GAN. Other generative architectures, such as Normalizing Flows and Diffusion Models, have also been used in MTSAD despite not directly modeling temporal or spatial dependencies. Normalizing Flows (e.g. MAF and Real NVP) transform a simple base distribution (e.g. Gaussian) into a complex target distribution via invertible mappings, allowing efficient density estimation [Kobyzev2021]. In MTSAD, anomalies are assumed to occur in low-density regions [GANF, MTGFlow]. Diffusion Models (e.g. DDPM) instead corrupt inputs with noise and learn to reverse this process for reconstruction [ImDiffusion, DDTAD]. In practice, these generative models are often combined with spatio-temporal architectures to better capture the complex MTS patterns.
@@ 4.4.4 Nested Models
In Nested models (Fig. 9(b)), typically, a temporal model is embedded within an AE a GAN-based model. For instance, several studies have implemented LSTM-AE architectures, where the encoder and decoder structure of an AE is composed of LSTM layers [Kieu2019, Nguyen2021]. As another example, the OmniAnomaly uses GRU incorporated into a VAE framework to create a stochastic RNN to learn latent representations and capture temporal dependencies [OmniAnomaly]. Additionally, DAEMON (adversarial Autoencoder anomaly detection interpretation) uses adversarial training in an AE-based model, where encoder and decoder networks consist of 1D-CNN, which apply filters that slide along the temporal dimension, to capture the temporal dependencies [DAEMON]. An example of a GAN-based model is MAD-GAN (Multivariate Anomaly Detection with GAN), which employs an LSTM network as the base model for both the Generator and Discriminator. MAD-GAN considers the entire dimension set simultaneously in order to capture the latent interactions among the different dimensions and uses LSTM to capture temporal dependency [MAD-GAN].
@@ 4.4.5 Parallel Models
Parallel models (Fig. 9(c)) handle both dependencies simultaneously using a two-branch architecture, one dedicated to temporal dependencies and the other to spatial dependencies. Typically, a temporal model is employed in one branch and a spatial model in the other. For example, the STADN (AD Network using Spatial and Temporal information) models spatial dependencies using a GAT and temporal dependencies using an LSTM. The two feature representations are then concatenated and passed through stacked fully connected layers for prediction [Tian2023]. Similarly, ESTAD (Enhanced Spatio-Temporal constraints end-to-end network for AD) employs a Transformer encoder to extract temporal features in parallel with a GNN model that learns spatial dependencies. It further introduces constraint tasks in both temporal and spatial domains, such as an autoregressive temporal objective and graph contrastive learning, to enhance spatio-temporal representations [ESTAD].
@@ 4.4.6 Other Models
The dominance of spatio-temporal models reflects the inherent complexity of MTS data, where anomalies often manifest in both temporal and spatial dependencies simultaneously. Temporal models are a natural choice when spatial dependencies are weak, while Spatial models are more appropriate when anomalies arise from disruptions in inter-sensor relationships. Among spatio-temporal configurations, Sequential models are appropriate when one dependency type is less critical compared to the other, Parallel models when both carry equal importance, and Nested models when temporal and spatial dependencies are tightly coupled and best captured within a unified architecture. Finally, Other models represent architectures that combine elements from multiple configurations in hybrid or unconventional ways. At this stage, the proposed taxonomy already provides a structured foundation that can be naturally extended as new architectural patterns consolidate in the literature.
@@ 5.1 Overview until 2024
Before 2019, MTSAD research was dominated by traditional ML methods, such as clustering, Support Vector Machines and tree-based models, which were valued for their interpretability and their ability to capture anomaly patterns across multiple variables. Although such works were identified in the survey, they were not included in the review since the focus lies on DL-based approaches. Readers interested in ML-based strategies are referred to existing surveys [BlzquezGarca2021, Paparrizos2025]. From 2019 onwards, a clear shift towards DL-based models can be observed. The first wave of contributions, particularly between 2019 and 2022, introduced disruptive architectural innovations, such as OmniAnomaly [OmniAnomaly], MAD-GAN [MAD-GAN], MSCRED [MSCRED], USAD [USAD], MTAD-GAT [MTAD-GAT], GDN [GDN], TranAD [TranAD], GTA [GTA] and Anomaly Transformer [AnomalyTransformer]. More recent works, published between 2023 and 2024, tend to focus on incremental improvements, either by fine-tuning existing architectures, modifying loss functions, or refining input and output representations. Nevertheless, notable exceptions remain, with models such as CAE-M [CAE-M] and DCdetector [DCdetector] introducing novel architectural concepts after 2022. Table 3 presents the categorization of these representative methodologies across the dimensions of the proposed taxonomy, illustrating their different architectural and methodological choices and enabling a structured comparison of the approaches within the taxonomy.
Table 3: Classification of a set of representative MTSAD methodologies on the dimensions of the proposed taxonomy.
Part Dimension
OmniAnomaly
MAD-GAN
MSCRED
USAD
MTAD-GAT
GDN
TranAD
GTA
Anomaly
Transformer
CAE-M
DCdetector
Input Preprocessing N N Y N N N N N N N N
Format W W I W W W W W W W W
Content Time Time Others Time Time Time Time Time Time Time Time
Output Type Rec Rec Rec Rec Hybrid Pred Rec Pred Rec Pred Rep
Optimal
Threshold
Y n.e. Y Y Y Y Y Y Y Y Y
Time-variyng
Threshold
N N N N N N Y N N N N
Granularity Point Point Subseq Point Point Point Point Point Point Subseq Point
Model Task Reg Reg Reg Reg Reg Reg Reg Reg Reg Reg Reg
Loss
Function
ELBO Adv SumE SumE Multi SumE SumE SumE Multi Multi Con
Feature
Extraction
N N N N ST S N ST N S T
T/S
Dependency
Nested Nested Nested S Others $Others S\rightarrowT T$
N: No; Y: Yes; W: Window; I: Image; Rec: Reconstruction; Pred: Prediction; Rep: Representation; Subseq: Subsequence; Reg: Regression; Adv: Adversarial; Multi: Multi-objective; Con: Contrastive; T: Temporal; S: Spatial; ST: Spatio-Temporal
@@
The categorization of 409 articles highlights several trends in the field for 2019–2024. Most studies rely on raw time series organized into sliding windows while preserving the temporal order, although there is a growing tendency to represent time series as images to capture spatial correlations. Regression-based settings dominate, and distance-based losses, such as MSE, remain the most widely adopted due to their simplicity, even though hybrid loss functions have been increasingly explored to capture more complex patterns. In temporal/spatial dependency modeling, sequential, nested, and other complex structures prevail, with nested and heterogeneous architectures being particularly common as these approaches jointly capture temporal and spatial dependencies. Moreover, parallel structures are gaining visibility due to their scalable nature, enabling simultaneous modeling of multiple dependencies and improving computational efficiency. Dedicated feature extraction modules are relatively rare, but when present, they typically focus on extracting spatio-temporal features. With respect to the output, anomaly scores are most often computed through reconstruction errors, although hybrid strategies that combine reconstruction errors with prediction errors or latent space deviations are gaining popularity. Detection continues to be primarily carried out at the point level, with thresholds that are generally optimal but time-invariant. The choice of architecture has also evolved, with Transformers and GNN becoming increasingly prominent. Transformers are particularly effective for temporal modeling, while GNN are widely adopted to capture spatial dependencies. More recent directions include the exploration of generative models, such as Normalizing Flows and Diffusion models, as well as contrastive learning strategies. The application of Large Language Models (LLM) to MTSAD is still at an early stage, but their adoption is increasing, reflecting the broader impact of Transformer-based architectures in time series research [GPT4TS, MOMENT, DualLMAD].
@@
The dimensions of the proposed taxonomy are independent, and each article is classified into a single category per dimension. Nevertheless, certain patterns may emerge across dimensions. The most frequent combination includes No Preprocessing, Window Format, Time Content, Reconstruction Type, an Optimal but Time-invariant Threshold, Point Granularity, Regression Task, ELBO Loss Function, Nested T/S Dependency, and No Feature Extraction module. This profile is largely driven by the popularity of LSTM-VAE [LSTM-VAE], one of the earliest DL approaches for MTSAD, which established reconstruction-based AD with uncertainty estimation as a dominant paradigm, with subsequent works building upon and extending its core methodology, as exemplified by OmniAnomaly [OmniAnomaly], which further incorporated GRU and NF to improve latent space modeling. However, this setup occurs in only 18 out of the 409 categorized papers, highlighting the heterogeneity of approaches. Since the Loss Function, Optimal Threshold and Time-varying Threshold dimensions highly contributed to this variability, they were ignored for subsequent analyses. When ignoring these three dimensions, the number of papers matching this combination rises to 79 out of 409 (i.e. 19.32%). This joint proportion is larger than the product of proportions over each dimension, which correponds to the case of categories being chosen independently ($\sim$5%). This indicates a positive association such that that the presence of a category in one dimension increases the likelihood of a category in the remaining dimensions.
@@
The proposed taxonomy is suitable to classify emerging paradigms applied to MTSAD, including foundation models such as TimesNet [TimesNet], LLM-based approaches [DualLMAD], and distributed approaches, such as federated learning-based methods [FDVL-DCN]. These models still rely on the same core design decisions captured by the taxonomy dimensions. As an example, DualLMAD [DualLMAD] employs two pre-trained GPT-2 models in a parallel fashion to simultaneously capture T/S dependencies within the time series. The model reconstructs the input window and computes the anomaly score as reconstruction error, using an optimal and time-varying threshold to obtain the final classification. While these approaches differ in their training paradigm, whether through large-scale pre-training or federated aggregation, their underlying detection methodology remains fully compatible with the proposed framework, confirming that it can accommodate future developments in MTSAD.
@@ 5.2 Some developments in 2025
Input-related dimensions (Pre-processing, Format and Content) show continuity with earlier years with some broadening. Most recent contributions continue to rely on Window-based formats for modeling subsequences. As examples, MLAD adopt fixed-length sliding windows to capture localized temporal dependencies, while TRLP introduce multi-resolution pyramid windows that model dynamics at multiple temporal scales. Together, these papers exemplify how contemporary research refines rather than replaces traditional window-based representations in MTSAD. At the same time, several studies move beyond purely temporal segmentation by integrating Window+Image representations or frequency-transformed inputs. For instance, FDTAD incorporate frequency-domain features alongside temporal windows, and FreqWave-TranDuD make use of wavelet-based frequency transformations to enhance multiscale sensitivity. These trends consolidate the Preprocessing Input dimension within the taxonomy and highlight a broader shift toward hybrid representations that merge temporal, spatial and frequency-domain information. Window-based formats remain dominant as they are compatible with most DL architectures and support the detection of Point, Contextual and Collective anomalies. The move toward hybrid content reflects the understanding that temporal information alone may be insufficient to capture all anomaly types, with frequency and spatial information providing complementary signals that improve detection.
@@
At the Output level, the categories of Reconstruction and Prediction remain central, but 2025 research expands the use of Representation-based deviations and Hybrid anomaly scores. AD2T (Association Discrepancy Dual-decoder Transformer) combines reconstruction, prediction and association discrepancy in a dual-decoder framework [AD2T], while TRLP (Tranformer-based Reconstruction-Limited Probability) introduces probabilistic scoring based on representation uncertainty [TRLP]. Furthermore, adaptive thresholding mechanisms gained traction e.g. Mixer-transformer employs time-varying thresholds to reduce false positives [Mixer-transformer], while boundary-aware ensembles generate thresholds based on estimated uncertainty [He2025]. Together, these papers reinforce that Hybrid type outputs are gaining ground as they exploit the complementary strengths of reconstruction, prediction and representation-based signals, improving the detection of diverse and complex anomaly patterns. Moreover, time-varying thresholds allow models to adapt to the non-stationary nature of real-world data. For Granularity, most contributions still detect anomalies at the Point level. For example, methodologies such as MLAD (Multi-task Learning Anomaly Detection) and MGDRF (Multi-Grain Dynamic Receptive Field) explicitly model multi-scale subsequences, yet they assign anomaly scores to each time point by considering its temporal context [MLAD, MDGRF]. These developments indicate that point granularity anomaly detection remains a natural choice for most DL-based MTSAD methodologies in 2025, as it provides the finest temporal resolution for anomaly localization and is directly compatible with window-based input formats.
The Model part is where 2025 contributions show the greatest methodological innovation. In the Task dimension, explicit Classification formulations remain uncommon and Regression continues to dominate, with models learning to identify deviations as anomalies [MLAD, TRLP, DiffANT, TiTAD, TSAD]. Moreover, several 2025 approaches integrate prediction with other representation learning (e.g. contrastive, variational or adversarial) to enhance the discriminative capacity of latent features [RGAnomaly, DyGCL, IGANomaly, AVAE]. As a consequence, the Loss Function dimension shows some relevant developments. Beyond the classical reconstruction losses (SumE, ELBO), the 2025 research incorporates more frequent Contrastive objectives (CARLA [CARLA], DMAP-DDCL [DMAP-DDCL]), Adversarial losses (EH-GAM-EGAN [EH-GAM-EGAN], Quantum GAN [Quantum_GAN]) and multi-objective criteria that merge reconstruction, prediction and representation learning (AD2T [AD2T], STAMP [STAMP]). These advances consolidate the Multi-objective category as a mainstream choice for the Loss Function dimension. Overall, this reflects a progression from simple error minimization toward richer training objectives, allowing models to capture more discriminative representations of normal behavior and thus improve sensitivity to subtle and complex anomalies.